Circular No. 6954

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 M4 (LINEAR) Parabolic orbital elements from 16 observations June 25–27:

T = 1997 Dec. 2.644 TT  ω = 102.^{o}022
Ω = 93.080 }2000.0
q = 2.47969 AU  i = 154.368
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 26 19^{h}44.^{m}37 −10^{o}12.^{′}9 2.349 3.302 155.^{o}7 7.^{o}3 16.5
July 1 19 31.37 −11 15.2 2.351 3.336 162.8 5.2 16.6
6 19 18.25 −12 17.0 2.367 3.370 168.7 3.4 16.6
11 19 05.22 −13 17.2 2.397 3.404 170.4 2.8 16.7
16 18 52.54 −14 14.6 2.442 3.438 166.5 4.0 16.8
21 18 40.40 −15 08.5 2.500 3.473 160.2 5.7 16.9
26 18 28.98 −15 58.6 2.572 3.508 153.2 7.5 17.0

SUPERNOVA 1998cp IN MCG +9-21-11 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag 18.0) on an unfiltered CCD frame taken with the 0.35-m Tenagra I automated supernova patrol telescope on June 21.27 UT. The candidate is located at α = 12^{h}31^{m}52^{s}.3, δ = +52^{o}23′39′′ (equinox 2000.0), which is 3′′ east and 24′′ south of MCG +9-21-11 (CGCG 270-008). The object is also present on frames taken on June 22.29 and 27.25, fading by ∼ 0.5 mag apparent by the latter occasion. The limiting magnitude of the exposures was ∼ 19. The object does not appear on either the first or the second Palomar Sky Survey scans, and it is also absent from the USNO-A1.0 catalogue.

METEORS I. Sato, National Astronomical Observatory, Tokyo, reports that sev-

eral observers in Japan observed a sudden meteor shower around June 27.6 UT. K. Usuki (Aizu, Fukushima) reported seeing 40–50 meteors per hour at the peak, even through 80–90-percent cloud cover; the radiant was between η UMa and α Boo. Y. Yabu (Omihachiman, Shiga) reported that an (unknown) observer in Okinawa saw “a couple” of meteors per minute, the radiant being in Corona Borealis. Y. Kushida (Yatsugatake, Yamanashi) recorded numerous meteors by VHF radio during June 27.58–27.67 UT. T. Hamane (Misato Astronomical Observatory) reported many bright meteors with trains.

1998 June 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden